OpenStudy (anonymous):

What is the equation to finding the number of sides a regular polygon has by its interior angles?

OpenStudy (anonymous):

are the polygons regular?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

just do the total angle sum divided by one interior angle size so 360 / 90 = 4 (square) 180 / 60 = 3 (triangle) 540 / 108 = 5 (pentagon) etc

OpenStudy (anonymous):

oh duh ha i feel stupid now, thank you. so 9*140=1,260

OpenStudy (anonymous):

the general rules are: n=number of sides Sum of Interior Angles = (n-2) × 180° Each Angle (of a Regular Polygon) = (n-2) × 180° / n